OpenStudy (anonymous):

PLEASE HELP! Express sin36degrees - sin32degrees as a product. a. -2cos34degrees sin2degrees b. 2cos34degrees sin2degrees c. 2sin34degrees cos2degrees d. -2sin34degrees cos2degrees

OpenStudy (anonymous):

Oh wait, I got b. Can you please clarify?

OpenStudy (anonymous):

\[\large \sin(C) - \sin(D) = 2\cos(\frac{C + D}{2}) \sin(\frac{C - D}{2})\]

OpenStudy (anonymous):

Well Done...

OpenStudy (anonymous):

Thanks!